The graph of the parametric equations is called a cycloid.

x=θ-sinθ and y=1-cosθ for 0≤θ≤2π

(a) find dy/dx

(b) find an equation of the tangent to the cycloid at the point where θ=π/3

(c) at what point is the tangent horizontal?

(d) graph the cycloid and the tangent lines in parts (b) and (c)


1
Expert's answer
2021-11-17T11:26:42-0500

a)

"dy/dx=\frac{dy}{d\theta}\frac{d\theta}{dx}"


"dy/d\theta=sin\theta"

"dx/d\theta=1-cos\theta"


"dy/dx=\frac{sin\theta}{1-cos\theta}"


b)

equation of the tangent:

"y-y_0=f'(x_0)(x-x_0)"


then, θ=π/3:

"1-cosθ-(1-cos(\pi/3))=\frac{sin(\pi/3)}{1-cos(\pi/3)}(θ-sinθ-(\pi/3-sin(\pi/3)))"

"1/2-cosθ=\sqrt3(θ-sinθ-\pi/3+\sqrt3/2)"

"cosθ+\sqrt3(θ-sinθ)=\pi/\sqrt3-1"


c)

the tangent horizontal if "f'(x)=0"

so,

"\frac{sin\theta}{1-cos\theta}=0"


"sin\theta=0"

"cos\theta\neq 1"

"\theta=\pi"

"y(\pi)=2,x(\pi)=\pi"


d)